Determine the final temperature of the liquid when you mix:

m1 = 1.00 kg water at t1 = 10.0 °C and m2 = 2.00 kg water at t2 = 40.0 °C.

Assume that no heat is lost to the environment.

FORMULAS:

Q1 + Q2 + .... + = 0

q= mcЛТ
Answer #1

Let the final temperature is T Q. + Q2 = 0 mic(T-T.) + M2 CLT-92)=0 ICT-10) 2 (T-40)=0 T-10 + 2 T-80 - 0 T=30°Clans